# A Comprehensive Survey of Foundation Models in Medicine

### v1

@@ Abstract
Foundation models (FMs) are large-scale deep-learning models trained on extensive datasets using self-supervised techniques. These models serve as a base for various downstream tasks, including healthcare. FMs have been adopted with great success across various domains within healthcare, including natural language processing (NLP), computer vision, graph learning, biology, and omics. Existing healthcare-based surveys have not yet included all of these domains. Therefore, this survey provides a comprehensive overview of FMs in healthcare. We focus on the history, learning strategies, flagship models, applications, and challenges of FMs. We explore how FMs such as the BERT and GPT families are reshaping various healthcare domains, including clinical large language models, medical image analysis, and omics data. Furthermore, we provide a detailed taxonomy of healthcare applications facilitated by FMs, such as clinical NLP, medical computer vision, graph learning, and other biology-related tasks. Despite the promising opportunities FMs provide, they also have several associated challenges, which are explained in detail. We also outline potential future directions to provide researchers and practitioners with insights into the potential and limitations of FMs in healthcare to advance their deployment and mitigate associated risks.
1 Introduction
1.1 History
@@
Artificial intelligence (AI) has achieved remarkable milestones in recent decades. This progress is mainly attributed to the enhanced availability of large-scale data, computational resources, and intelligent algorithm design [1]. AI-based techniques have been used for decades within healthcare [2]. For example, computer-assisted surgery began in the early 1980s, which marked the start of medical AI [3]. In 1989, a small convolutional neural network (CNN) model demonstrated promising performance by using a backpropagation algorithm to recognize handwritten digits [4]. However, the application of AI in medicine was not facilitated until the AI spring, which was initiated by the emergence of the AlexNet model [5]. The significant momentum in the development of AI models was demonstrated after the arrival of the AlexNet model. This was trained on the most prominent training dataset called ImageNet [6]. Subsequently, other state-of-the-art deep learning (DL) models have been developed to surpass AlexNet [5] through the ImageNet Large Scale Visual Recognition Challenge (ILSVRC) [6]. Following the CNN, Recurrent neural networks (RNN) [7] were proposed for sequential data processing that showed promising performance in various domains including time series analysis, natural language processing (NLP), and speech recognition [8]. Another significant contribution came with the introduction of generative adversarial networks (GAN) [9] for generating synthetic data [9, 10].
@@
Transformers, introduced in 2017, formed a crucial foundation of the most powerful deep learning models available today. Transformers introduced a self-attention mechanism to capture sequences in parallel and addressed the limitations of computational and sequential RNNs. Various state-of-the-art models have been developed based on the success of transformers. Transformer models have also played a key role in driving the creation of foundation models (FM) [11]. FMs are large-scale pre-trained neural networks trained on large amounts of data that can be fine-tuned to perform a wide range of downstream tasks. The key enablers of FMs include training data, base model, transfer learning, and scale [11]. FMs adopt a self-supervised learning (SSL) approach, unlike traditional DL models that require a large amount of data. In this approach, the pre-training task is autonomously generated from unlabeled data. SSL evolved from word embeddings that provided context-independent vectors for words [12, 13] to autoregressive language modeling that predicts the next word based on previous words [14, 15].
@@
SSL led to one of the first FMs known as BERT [16]. A powerful transformer architecture characterized by its bidirectional sentence encoders and scalability extending to larger architectures and datasets. BERT became the standard model in 2019. Many subsequent models were developed based on BERT architectures, including RoBERTa [17], BART [18], and T5 [19]. These models have become prevalent in the development of state-of-the-art NLP models. These NLP models are termed large language models (LLM). Despite being available since late 2018, these models gained broader recognition across various domains following the introduction of ChatGPT in late 2022111https://chat.openai.com/. Furthermore, they were soon generalized to different fields, such as text [20], images [21], videos [21], speech [22], tabular data [23], protein sequences [24], organic molecules [25], and reinforcement learning [26].
@@
Healthcare constitutes a substantial economic sector representing 17% of the United States GDP [11]. Many specialized models have been developed for various healthcare tasks [27]. However, they are based on a limited amount of labeled data and face challenges when generalized to different tasks. It remains challenging to adopt general-purpose FMs in the healthcare domain due to the complex structures and limited publicly available medical data. Therefore, several works have been carried out to define FMs specifically for medical data. These models are task-agnostic and can be generalized across different tasks within the medical domain. For instance, the GatorTronGPT [28] model was built on the GPT-3 [29] architecture and includes 20 billion parameters. It was trained on 277 billion words with 82 billion words of clinical text which were obtained from 2 million patients. GatorTronGPT exhibits strong performance in various medical tasks such as named entity recognition (NER), biomedical question answering (QA), biomedical relation extraction, natural language inference (NLI), and semantic textual similarity. Similarly, models such as MedSAM [30] trained on 1.5 million medical images, are used for medical image segmentation and surpass other models [31, 32]. FMs designed for other tasks within the medical domain are described in detail in Section 5.
1.2 Research Objectives and Outline
@@
This survey paper aims to provide a comprehensive review of FMs in the healthcare domain. We first provide an overview of the historical progression of artificial intelligence (AI) and deep learning (DL) models followed by the development of transformers-based models (shown in Figure 1). The survey is structured to provide a comprehensive study of FMs, with Section 2 explaining the background of FMs, followed by an extensive overview of FMs in Section 3. The flagship FMs such as the BERT family, GPT models, CLIP, stable diffusion, and industrially scaled pre-trained models are explained in Section 4. The core FMs specifically designed for medical applications are presented in Section 5. The applications of FMs in the healthcare domain are presented in Section 6. We present multiple applications including clinical NLP (Section 6.1), medical computer vision (Section 6.2), healthcare graph learning (Section 6.3), biology and omics (Section 6.4), and others (Section 6.5). In Section 7, we explain the available opportunities associated with FMs, including controllability, adaptability, and applicability. The challenges associated with FMs, such as cost, interpretability, validation, and scale, are thoroughly discussed in Section 8. A discussion and potential future work are presented in Section 9 and a conclusion is provided in Section 10.
Figure 1: Outline of the proposed work.
1.3 Scope
@@
FMs have been widely adopted across multiple domains. However, a detailed taxonomy covering all aspects of FMs within the medical domain is still missing. Existing surveys have concentrated on general-purpose FMs [11, 33, 34, 35, 36], whereas this survey focuses on the healthcare domain. Furthermore, we describe the major flagship models that led to the development of medical-based FMs (Section 4).
1.4 Existing surveys
@@
In 2023, the rapid advances in FMs led to the publication of multiple survey articles (Table 1). However, the majority of these articles do not focus on healthcare. Some of the articles only focus on LLMs on [37], which excluded FMs for other modalities such as diffusion models [38]. Furthermore, existing articles focused on healthcare lack a detailed taxonomy of FMs that includes all of its associated aspects, such as LLMs, vision, protein, audio, and graph FMs.
Table 1: Comparison with existing survey.
@@ 1.5 Contribution
This survey provides a comprehensive analysis of recent works published until early 2024. It encompasses the evolutionary journey of FMs in the healthcare sector, providing insights into their historical development and categorization. The survey explores the applications of FMs in various healthcare domains including clinical NLP, medical computer vision, biology and omics, and other relevant areas. Furthermore, we highlight the major challenges associated with FMs and outline future directions associated with FMs in healthcare.
2 Background
2.1 Transformer
A Transformer [44] is a neural network architecture initially developed for sequence modeling. It addressed the inefficient computation and poor performance of RNNs caused by the sequential processing of the information. Due to the attention mechanism in transformers, sequences can be processed in parallel, which is faster and more efficient during training. The attention mechanism in a transformer is computed by taking a weighted sum of the representations of all tokens in a sequence, where the weights are determined by the similarity between the tokens. Since there is no convolution or recurrence, positional encoding is used to represent the relative positions of tokens in a sequence. The positional encoding vectors are added to the embedding vectors for the tokens at the input to the self-attention mechanism. This helps with learning long-range dependencies in a sequence by attending to the positional encoding vectors of the tokens. Transformers have led to the development of state-of-the-art models including BERT [16], GPT [15][45], and other models [35]. Transformers were originally developed for NLP tasks, but they have since been extended to other domains, including computer vision. Vision transformer (ViT), first proposed in 2021 [46] are a type of transformer specifically designed for computer vision tasks. It converts an image into a sequence of patches and then calculates meaningful embedding using a self-attention mechanism. The positions of patch embeddings are retained using positional encoding.
2.2 Attention
Attention mechanism in deep learning originated from the human visual attention that selectively focuses on one or more aspects of information within a vast array of available data [47]. For instance, Itti et al. [48] proposed visual attention, inspired by the primate visual system, to identify salient regions. The attention mechanism has been used mainly with RNNs or CNNs [49] for multiple tasks [47]. However, transformers rely solely on attention, unlike CNNs and RNNs which use attention in part of the model. The attention mechanism in transformers computes the similarity between tokens in a sequence. The attention function maps a query and a set of key-value pairs to an output, where the output is a weighted sum of the values, with the weights determined by the similarity between the query and each key [44]. Multiple attention mechanisms have been proposed recently, including self-attention [44], multi-head attention [44], masked attention [50], and criss-cross attention mechanisms [51]. In self-attention, each element in the input sequence is associated with query, key, and value vectors. The model then computes attention weights by measuring the similarity between the query and key vectors for all elements in the sequence. These weights guide the model in taking a weighted sum of the value vectors for each element, effectively enabling the model to selectively focus on different parts of the input when processing each element. Multi-head attention learns multiple representations of the input sequence, each of which can attend to different aspects of the sequence. This is done by performing multiple self-attention operations in parallel, on different input representations. Cheng et al. [50] proposed Mask2Former which utilizes masked attention in the transformer decoder. Masked attention restricts attention to specific regions of interest, such as the foreground region of a predicted mask, during the calculation of attention scores. Criss-cross attention [51] efficiently captures full-image contextual information for tasks like semantic segmentation and object detection. It utilizes a novel Criss-Cross Network with a criss-cross attention module, allowing each pixel to gather contextual information from its surroundings.
2.3 Self-supervised learning (SSL)
The limitation of supervised learning is that it depends heavily on large amounts of labeled data, which is a time-consuming, expensive, and often infeasible task [52] [53]. Furthermore, since the data are annotated based on the labels due to correlation, it introduces generalization error [52] and is prone to adversarial attacks [53]. In contrast, self-supervised learning (SSL) can be used to train the model on large amounts of data without labels. SLL leverages the inherent structure or context within the data itself to create meaningful representations. SSL can be divided into three main categories: contrastive, generative, and adversarial learning [53]. Contrastive learning algorithms train models to distinguish between similar and dissimilar data points based on similarity within the representation space. Generative learning algorithms train models to generate realistic data. Adversarial learning is a hybrid generative-contrastive learning approach that improves representations by reconstructing the original data distribution rather than individual data samples, using discriminative loss functions as objectives.
2.4 Human feedback reinforcement learning
Reinforcement learning (RL) is an ML-based algorithm in which an agent interacts with an environment and learns to make a sequence of decisions to maximize a cumulative reward [54]. It can be formulated using a Markov Decision Process (MDP), mathematically, $\langle S,A,T,R,\gamma\rangle$ [54, 55], where $S$ represents multiple states in an environment and selects actions ($A$). The transition probability ($T$) specifies how the agent moves from one state ($s$) to another ($s^{\prime}$), and a reward function ($R$) provides feedback on the actions taken. The discount factor ($\gamma$) is a critical parameter that influences the relative importance of future discounted rewards in the decision-making process. The agent’s goal is to find a policy ($\pi:S\rightarrow A$) that specifies how to select actions in each state to maximize the expected reward. Although RL learns complex behavior with predefined reward functions, however, it can be challenging to express some behaviors as rewards [55]. Therefore, human feedback RL (HFRL) in which the agent learns about their action from feedback from a human trainer. The human feedback has resulted in significant performance improvements [56].
2.5 CLIP
Traditional DL algorithms have demonstrated promising performance in multiple computer vision tasks. However, they rely heavily on extensive labeled training data, which is both labor-intensive and expensive to obtain. Additionally, these algorithms struggle to generalize beyond the specific tasks they are trained for [57, 58, 59]. These challenges are addressed by introducing CLIP [57] which is a novel neural network-based architecture that leverages natural language-supervised vision data readily available on the internet, effectively reducing the dependence on expensive labeled datasets while enabling broader generalization.
CLIP is a robust neural network architecture with zero-shot capabilities [60] that achieves various classification benchmarks using natural language instructions, without direct optimization for each specific benchmark [60]. It is a simplified variant of ConVIRT [61] and is trained on a dataset called WebImageText (WIT), comprising 400 million pairs of images and corresponding text collected from the internet. Multiple works such as [61, 62, 63, 64]utilized text-image representations, however, these methods operate under supervised, weakly-supervised, self-supervised, or unsupervised learning paradigms. In contrast, CLIP utilizes natural language as its training signal. The key advantage of leveraging natural language supervision for learning is its scalability, unlike traditional crowd-sourced labeling which requires classic ML-compatible annotations. CLIP can passively learn from the abundant textual data available on the internet, establishing connections between visual representations and language, which enables flexible zero-shot transfer to a wide range of tasks
CLIP employs separate transformer encoders for text and image inputs. The visual encoder is based on the ViT, while the text encoder utilizes the GPT-2 transformer. To enhance efficiency, CLIP employs a contrastive objective, predicting which text corresponds to a given image. It jointly trains an image and a text encoder to maximize the similarity between embeddings of paired (image, text) examples while minimizing similarity with other combinations. This process results in a multi-modal embedding space where images and their associated texts are closely related. The existing variations of CLIP-based models include CRIS [65], CLIPseg [66], FLIP [67], SLIP [68], MaskCLIP [69], and GLIP [70].

### v2

@@ Abstract
Foundation models (FMs) are large-scale deep-learning models that are developed using large datasets and self-supervised learning methods. These models serve as a base for different downstream tasks, including healthcare. FMs have been adopted with great success across various domains within healthcare. Existing healthcare-based surveys have not yet included all of these domains. Therefore, we provide a detailed survey of FMs in healthcare. We focus on the history, learning strategies, flagship models, applications, and challenges of FMs. We explore how FMs such as the BERT and GPT families are reshaping various healthcare domains, including clinical large language models, medical image analysis, and omics. Furthermore, we provide a detailed taxonomy of healthcare applications facilitated by FMs, such as clinical NLP, medical computer vision, graph learning, and other biology-related tasks. Despite the promising opportunities FMs provide, they also have several associated challenges, which are explained in detail. We also outline open research issues and potential lessons learned to provide researchers and practitioners with insights into the capabilities of FMs in healthcare to advance their deployment and mitigate associated risks.
Index Terms:
Foundation Models, large-scale healthcare models, Natural language generation, large language models, Contrastive Learning, diffusion models, self-supervised learning.
I Introduction
I-A History
@@
Artificial intelligence (AI) has reached significant milestones in recent decades. This progress is mainly attributed to the enhanced availability of large-scale data, computational resources, and intelligent algorithm design [1]. AI-based techniques have been used for decades within healthcare [2]. For example, computer-assisted surgery began in the early 1980s, which marked the start of medical AI [3]. In 1989, a small convolutional neural network (CNN) model demonstrated promising performance by using a backpropagation algorithm to recognize handwritten digits [4]. However, the application of AI in medicine was not facilitated until the AI spring, which was initiated by the emergence of the AlexNet model [5]. The significant momentum in the evolution of AI models was demonstrated after the arrival of the AlexNet model. Subsequently, models such as VGGNet and ResNet further advanced the field of image classification, object detection, and segmentation. Following CNN, Recurrent neural networks (RNN) [6] were proposed for sequential data processing that showed promising performance in various domains, including time series analysis, natural language processing (NLP), and speech recognition [7]. Another significant contribution came with the introduction of generative adversarial networks (GAN) [8] for generating synthetic data.
@@
Transformers, introduced in 2017, formed a crucial foundation of the most powerful deep learning models available today. Transformers introduced a self-attention mechanism to capture sequences in parallel and addressed the limitations of computational and sequential RNNs. Initially designed for NLP, transformers have been extended to other fields, including computer vision such as Vision Transformer (ViT), introduced in 2021 [9], which is tailored specifically for vision tasks. Transformer models have led to the development of foundation models (FM) [10], which are large-scale pre-trained models trained on vast amounts of data and fine-tuned to perform a wide array of downstream tasks. The key enablers of FMs include training data, base model, transfer learning, and scale [10]. FMs adopt a self-supervised learning (SSL) approach, unlike traditional DL models that are dependent on large amounts of data. In this approach, the pre-training task is autonomously generated from unlabeled data. SSL evolved from word embeddings that provided context-independent vectors for words to autoregressive language modeling, which predicts the subsequent word based on previous words [11, 12].
@@
SSL led to one of the first FMs known as BERT [13], a powerful transformer architecture characterized by its bidirectional sentence encoders and scalability extending to larger architectures and datasets. BERT became the standard model in 2019. Many subsequent models were developed based on BERT architectures, including RoBERTa [14], BART [15], and T5 [16]. These models have become prevalent in the development of state-of-the-art (SOTA) NLP models. These NLP models are termed large language models (LLM). Despite being available since late 2018, these models gained broader recognition across various domains following the introduction of ChatGPT in late 2022111https://chat.openai.com/.
Similarly, vision-based models such as ViT, Swin Transformers [17], and Data-efficient Image Transformer [18] have demonstrated SOTA performance in visual recognition tasks. Once pre-trained, they can be fine-tuned for various healthcare applications. Furthermore, these models were soon generalized to different fields, such as text [19], videos [20], speech [21], tabular data [22], protein sequences [23], organic molecules [24], and reinforcement learning [25].
@@
Healthcare constitutes a substantial economic sector representing 17% of the United States GDP [10]. Many specialized models have been developed for various healthcare tasks [26]. However, they are based on limited labeled data and face challenges when generalized to different tasks. It remains challenging to adopt general-purpose FMs in the healthcare domain due to the complex structures and limited publicly available medical data. Therefore, several works have been carried out to define FMs specifically for medical data. These models are task-agnostic and can be generalized across different tasks within the medical domain. For instance, the GatorTronGPT [27] model was built on the GPT-3 [28] architecture and includes 20 billion parameters. It was trained on 277 billion words with 82 billion words of clinical text, which were obtained from 2 million patients. GatorTronGPT exhibits strong performance in various medical tasks such as named entity recognition (NER), biomedical question answering (QA), biomedical relation extraction (RE), natural language inference (NLI), and semantic textual similarity. Similarly, models such as MedSAM [29] trained on 1.5 million medical images are used for medical image segmentation and surpass other models [30, 31]. FMs designed for other tasks within the medical domain are described in detail in Section II.
I-B Research Objectives and Outline
@@
This survey paper aims to comprehensively review FMs in the healthcare domain. First, we present an overview of the history of artificial intelligence (AI) models, followed by the development of transformers-based models (Figure 1). The survey is structured to provide a comprehensive study of FMs, explaining the background of FMs (Supplementary Materials Section I), followed by an extensive overview of FMs (Supplementary Materials Section II). The flagship FMs such as the BERT family, GPT models, CLIP, stable diffusion, and industrially scaled pre-trained models, are explained in Supplementary Materials Section III). The core FMs specifically designed for medical applications are presented in Section II. The applications of FMs in the healthcare domain are presented in Section III. We present multiple applications, including clinical NLP (Section III-A), medical computer vision (Section III-B), healthcare graph learning (Section III-C), biology and omics (Section III-D), and others (Section III-E). Section IV-B explains the available opportunities associated with FMs, including controllability, adaptability, and applicability. The challenges associated with FMs, such as cost, interpretability, validation, and scale, are thoroughly discussed in Section IV-C. We highlighted the open research question and potential lessons learned in Section IV-D, and a conclusion in Section V.
Figure 1: Outline of the survey paper. We present an overview of medical FMs, followed by their applications in healthcare. Additionally, we offer in-depth insights into the taxonomy of medical FMs, the opportunities they enable, the challenges they face, and open research questions and lessons learned.
I-C Scope
@@
FMs have been widely adopted across multiple domains. However, a detailed taxonomy covering all aspects of FMs within the medical domain is still missing. Existing surveys have concentrated on general-purpose FMs [10, 32, 33, 34, 35], whereas this survey focuses on the healthcare domain. Furthermore, we describe the major flagship models that led to the development of medical-based FMs (Supplementary Materials Section III).
I-D Existing surveys
@@
In 2023, the rapid advances in FMs led to the publication of multiple survey articles (Table I). However, the majority of these articles do not focus on healthcare. Some articles only focus on LLMs on [36], which excluded FMs for other modalities such as diffusion models [37]. Furthermore, existing articles focused on healthcare lack a detailed taxonomy of FMs that includes all of its associated aspects, such as LLMs, vision, protein, audio, and graph FMs.
TABLE I: Comparison with existing survey.
@@ I-E Contribution
This survey provides a detailed analysis of recent works published until early 2024. It encompasses the evolutionary journey of FMs in the healthcare sector, providing insights into their historical development and categorization. The survey explores the applications of FMs in various healthcare domains including clinical NLP, medical computer vision, biology and omics, and other relevant areas. Furthermore, we highlight the major challenges associated with FMs and outline future directions associated with FMs in healthcare.
II Foundation Models in the Medicine Domain
Multiple FMs have been developed for general purposes that can be adopted across different domains. The main architectures are encoder, decoder, and encoder-decoder (Supplementary Materials Section-II A). The distribution of the representative models depending on their learning architectures is shown in Figure 2. A significant number of FMs are based on architectures like BERT, GPT, and Stable Diffusion. Therefore, brief explanations of these models are provided in the Supplementary Materials (Section III).
Although multiple general-purpose FMs are developed, directly adopting them for healthcare could be challenging. For instance, models such as Word2Vec, ELMo, and BERT, primarily trained on general domain datasets, make their performance on biomedical texts uncertain due to significant differences in word distributions [44]. Similarly, employing vision-text pre-training models like CLIP poses challenges because of its extensive data requirements and the subtle, fine-grained differences in medical images compared to general images [45]. Furthermore, the diverse imaging modalities in medicine complicate the development of unified models, as seen with the Segment Anything Model (SAM), which is ineffective for 3D volumetric medical images due to its 2D architecture [46]. Therefore, there is a need for specialized FMs tailored to the unique demands of the healthcare domain. These emerging models have shown to be highly successful (Table LABEL:tab:Flagship). This section highlights FMs exclusively designed for healthcare applications that tackle challenges and harness opportunities within the healthcare domain.
II-A Clinical Large Language Models (CLLM)
BioBERT is one of the earliest domain-specific LLMs [44]. It is a biomedical language processing model that was trained on large-scale biomedical corpora. BioBERT demonstrated impressive performance across 15 datasets addressing three NLP tasks: NER, RE, and QA. Another model called BioMegatron [47] employs a BERT-style architecture with up to 1.2 billion parameters. It is trained on a larger domain corpus (PubMed abstracts and PMC full-text) that outperforms previous SOTA models on biomedical NLP benchmarks (NER, RE, QAs). GatorTron [48] is another SOTA large-scale clinical language model developed from scratch was designed for NLP tasks on EHRs. It has 8.9 billion parameters and is trained on a corpus of over 90 billion words. These data included substantial clinical text from the University of Florida (UF) Health, PubMed, Wikipedia, and MIMIC III [49]. It outperforms current biomedical and clinical transformers across five distinct clinical NLP tasks. ChatGPT is a general-purpose model trained on various datasets, including a portion of healthcare data. However, the primary focus is not limited to the healthcare domain. Inspired by ChatGPT, GatorTron [48] was further extended to the GatorTronGPT [27] model. This was built on the GPT-3 architecture with 20 billion parameters. The training corpus included 277 billion words with 195 billion words from the Pile dataset and 82 billion words from de-identified clinical text obtained from UF Health. GatorTronGPT exhibits a remarkable ability to generate synthetic data (GatorTronS) aligned with medical knowledge, as validated through various clinical tasks. Furthermore, the performance on different tasks including NER, QA, RE, NLI, and semantic textual similarity achieved promising performance. Moreover, it has been tested on the physicians’ Turing test and achieved comparable performance to humans. Singhal et al. [50] proposed Med-PaLM, evaluated on a well-curated MultiMedQA benchmark. The proposed model achieved promising performance; however, it was slightly inferior compared to clinicians. Another model built upon PaLM-2 (Med-PaLM-2) was fine-tuned on the MultiMedQA benchmark and achieved 19% better performance than Med-PaLM. The more recent MEDITRON model [51] is a large-scale medical language model with 70 billion parameters trained with 48.1 billion tokens from the medical domain. It has outperformed previous SOTA models including GPT-3.5 and MedPaLM.
II-B Medical Image Analysis
FMs are trained on large numbers of training images, enabling fine-tuning for diverse tasks. However, the patterns and features in medical images are different from those in natural images [52]. Multiple works have been done for robust domain-specific models for medical images. For instance, the Segment Anything Model (SAM) [31] was originally developed for general-purpose segmentation and trained on over 11 million images. It has been fine-tuned for medical image applications with great success shown by [53]. A medical-related FM for universal medical image segmentation called MedSAM was proposed by Ma et al. [29] which employs a vision transformer-based (ViT) image encoder for feature extraction and a mask decoder for segmentation results. MedSAM uses multiple online medical datasets with 1,570,263 images and text pairs for training. A detailed evaluation showed promising performance compared to other segmentation models, including SAM [31] and U-Net [30]. Following the success of MedSAM, numerous models have emerged for medical image segmentation [54, 46, 55]. The details of these segmentation tasks are explained in Section III-B3.
II-C CLIP-based FMs